OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $395. How many sophomores bought tickets? 55 75 90 95

OpenStudy (anonymous):

Use a system of equations; You know there are 150 students total. f+s=150 You know the price of the tickets and the fee per freshmen/sophomore. 2f+3s=395

Directrix (directrix):

f+s=150 2f+3s=395 ---------- Simultaneously solving 2 equations in 2 unknowns, f and s, by substitution f + s = 150 f = 150 - s 2(150 - s) + 3s = 395 300 - 2s + 3s = 395 300 + s = 395 s = 95 -----> number of sophomores f = 150 - 95 = 55 ---> number of freshmen